# Hidden in Plain Sight II: Characterizing the luminous companion to $\kappa$ Velorum with VLTI/GRAVITY

### v1

@@ 5 UVES Spectra
Figure 6 shows the UVES spectrum and the synthetic spectrum of the best-fitting B-star model and marks hydrogen Balmer, He I and metal lines. The synthetic spectrum generally reproduces the observed He I and metal lines, but less accurately matches the wings of the hydrogen Balmer lines, especially at $\lambda<4000$ Å. This is likely to represent an artifact of the continuum normalization procedure in areas of the spectrum dominated by the broad Balmer line wings. The Si II and O II lines between $4550\lesssim\lambda\lesssim 4700$ Å are also under predicted by the best matching synthetic template, which could suggest that $\kappa$ Vel is alpha-enhanced. This analysis also assumes that the spectrum is entirely dominated by the photometric primary. We evaluate this assumption in more detail in Section §6.
The spectrum is consistent with a B2V-IV star. In addition to having no He II lines, the Si II doublet at $\lambda$4129 is clearly visible, and the C II $\lambda$4267 line is strong. The spectrum is qualitatively similar to the B2V spectroscopic standard $\beta^{2}$ Sco (Negueruela24). The most notable differences between the best matching synthetic spectrum and the UVES spectrum of $\kappa$ Vel are the He I lines at $\lambda\lambda$4471, 4922. Figure 7 shows a zoom in on these absorption lines. In both cases, we observe an additional blended component in the blue wing of the primary absorption feature that is not represented by our synthetic spectra at any effective temperature, surface gravity, or metallicity. We consider two possibilities: (a) the additional component is actually the He I absorption line of the fainter B-star companion or (b) the line list we use with Synspec does not include this blended absorption line.
@@
While the structure in the blue wings of the He I lines at 4471Å and 4922Å are not predicted by our synthetic spectra, similar He I lines are found in other B and Be stars. For example, Negueruela24 show that the B3V star HD 178849 has structure in the He I $\lambda$4471 line that is not apparent in stars of the same spectral class with higher rotational velocities. Similarly, the spectrum of HD 58343, a Be2V star with $v\sin i=43$ km/s also has a similar $\lambda$4471 line profile (Chauville01) which we show in Figure 7 for comparison. The most likely explanation is that the blue component of the $\lambda$4471 comes from the forbidden transition [He I] at $\lambda$4469.96 (Daflon07). This is is less commonly observed because it requires a high-resolution spectrum and the star to have projected rotational velocity $v\sin i\lesssim 50$ km/s (see Figure 6 of Santos25). A forbidden [He I] line is also expected at $\lambda$4921 (Beauchamp98), accounting for the blue wing in the He I $\lambda$4922 profile seen in Figure 7.
Figure 6: High-resolution UVES spectrum (black) taken as part of the EDIBLES survey (Cox17). The red line shows the best fitting synthetic spectrum computed using TLUSTY model atmospheres and Synspec. Hydrogen, helium, and metal lines are marked using wavelengths from the NIST database (NIST_ASD).
Figure 7: UVES spectrum of $\kappa$ Vel showing the two He I absorption lines with asymmetric blue wings not predicted by the synthetic spectra. The $\lambda$4471 line of the Be2V star HD 58343 from Chauville01 is shown for comparison. We find that the apparent second component in the blue wings of these absorption lines is more likely from forbidden [He I] transitions rather than the stellar secondary.
6 Comparison to Evolutionary Tracks
@@
By combining the measured binary masses and the flux ratio from the VLTI GRAVITY observations, we can estimate the evolutionary state of the binary and the spectroscopic properties of the secondary. We download precomputed evolutionary tracks (Dotter16; Choi16) from MESA Isochrones and Stellar Tracks. For a given mass pair of $M_{1}$ and $M_{2}$ and metallicity, we interpolate their tracks onto the same age grid and calculate the $K$-band flux as a function of stellar age. We can then estimate the properties of both stars and the age of the system.
@@
The measured masses are strongly dependent on the assumed parallax (Figure 5). The parallax uncertainty is large ($0.3$ and $0.48$ mas for the original and updated measurement, respectively) and the parallax could be biased by the orbital motion by as much as 22% (Section §4). The effective temperature of the primary star measured from UVES can be used to place additional constraints on the masses of the stars. Figure 8 shows the evolution of the flux ratio of the binary system as a function of the effective temperature of the primary star assuming different parallaxes. We assume a metallicity $\rm{Fe/H}]=-0.3$ based on our template minimization results. The measured flux ratio, $\epsilon=0.285\pm 0.003$, implies dramatically different primary effective temperatures depending on the parallax and the derived primary mass. For example, assuming the original Hipparcos parallax $\varpi=6.05$ mas, the flux ratio measurement implies that the photometric primary is still about halfway through its main sequence lifetime and has an effective temperature of $T_{\rm{eff},1}>25$ kK, which is much hotter than the temperature we estimate from the UVES spectrum. Figure 8 shows that the effective temperature estimate from UVES and binary flux ratio prefer a parallax of $\varpi\approx 6.5$ mas. This is $\gtrsim 1\sigma$ larger than the original Hipparcos parallax measurement and $\gtrsim 3\sigma$ larger than the revised Hipparcos parallax measurement using the reported uncertainties, but within the broader range of parallaxes acceptable if the measurement has been biased by the binary motion.
@@
Assuming this value for the parallax, we can asses whether or not we expect to see any spectral features of the secondary in the UVES spectrum. The primary and secondary masses are $M_{1}=7.5\ M_{\odot}$ and $M_{2}=5.9\ M_{\odot}$, respectively. We use these masses and the orbital solution and predict that velocity difference between the primary and the secondary star at the time of the UVES observation is only $\Delta\rm{RV}=12.2$ km/s, which is not surprising given that the UVES observation was taken near conjunction. At a system age of $\sim 40$ Myr, which is when the $K$-band flux ratio matches the VLTI GRAVITY result, both stars have $T_{\rm{eff}}\approx 19{,}000$ K, so the flux ratio is roughly the same in the $K$-band and in the wavelength range of the UVES spectrum. We take the synthetic spectra from the TLUSTY BSTAR2006 grid described in Section §5 and select the closest matches to the primary and secondary star. We broaden these to match the UVES resolution and apply the expected RV shifts for each component. We also rotationally broaden the spectra of both components by 40 km/s. We combine the synthetic primary and secondary spectra, weighting them such that the flux ratio is $\epsilon=0.28$. We find that the differences between the spectrum of the photometric primary and the combined binary spectrum are at the $\lesssim 2\%$ level. It is therefore unsurprising that we do not detect any signatures of the secondary in the UVES spectrum.
@@ 7 Discussion and Conclusions
$\kappa$ Vel is one of the brightest stars in the Southern sky and was first detected as a single lined spectroscopic binary by Curtis1907. Here, we use VLTI GRAVITY to definitively determine the nature of the secondary star for the first time. To the best of our knowledge, Daszynska17 present the only previous evidence that the companion is a massive star and not a compact object. They used BRITE photometry to characterize $g$-mode oscillations and found a frequencies inconsistent with the spectral type and evolutionary state of the B2V-IV photometric primary. They instead suggest that the oscillations come from a fainter, less massive B star companion.
@@
We observed $\kappa$ Vel star as part of a broader effort to identify stars with compact object companions. Given the predicted large angular separation (Figure 1) and large companion mass (Figure 2), $\kappa$ Vel emerged as a promising candidate for VLTI GRAVITY observations. Although were were unambiguously able to identify the luminous secondary in the first VLTI observation, we obtained a total of nine epochs, and simultaneously fit the astrometry and RVs to model the binary orbit (Figure 3). From this model, we can predict the masses of the binary components (Figure 5 and Table 4). We find that the uncertainty in the mass measurement is primarily driven by the uncertainty in the Hipparcos parallax. It is also likely that the binary motion has biased the parallax measurement by as much as $\sim 22\%$.
@@
Given its spectral type and magnitude, $\kappa$ Vel has also been targeted by studies of diffuse interstellar bands. We use the publicly available UVES spectrum from Cox17 and B-star model atmospheres to estimate the effective temperature, surface gravity, and rotational velocity of the photometric primary (Figure 6). By combining the binary flux ratio measured from VLTI GRAVITY and the effective temperature of the photometric primary, we can estimate the distance and binary masses (Figure 8).
@@
We also use this model to estimate how the secondary star contributes to the combined spectrum. Even at RV quadrature, when the velocity difference between the two stars is largest, the synthetic binary system is expected to differ from the photometric primary at the few-percent level (Figure 9). While it may be possible to detect the subtle spectroscopic signatures of the secondary with multi-epoch high-resolution spectra and techniques like spectral disentangling (e.g., Shenar20; Seeburger24), we detect the luminous secondary with the first VLTI GRAVITY observation. Since the orbital period and ephemeris is known from the RVs, a single VLTI observation could be used to rule out a black hole with a total observing time of $\sim 1$ hour, including overheads. For comparison, the single UVES spectrum shown in Figure 6 had a total observing time of 23 min, and multiple spectra ($\gtrsim 5$) taken at different orbital phases are typically needed for spectral disentangling to be effective.
7.1 Other VLTI Targets
@@
X Per 5.9 3900.0 6.6 20.7 Be X-ray binary pulsar (Fortin23)
HD 45910 4.4 232.5 1.7 8.1 B2+KIII (Cowley64)
$\chi$ Lup 4.0 15.3 4.1 7.2 B9.5V+ A2V (LeBouquin13)
$\delta$ Gem 2.6 2238.6 353.1 6.0 Uncertain RV orbit (Abt74)
HD 187299 3.5 1901.0 7.4 5.9 G5I + B7V (Griffin79)
HD 50975 5.7 190.2 0.8 5.2 F8I + B2 (Sperauskas14)
@@
HD 8556 4.9 5895.0 240.3 4.1 F0 + F5 (Fletcher73)
71 Tau 4.0 5200.0 227.4 3.9 HST UV emission lines (Simon00)
HD 127208 5.1 24.6 1.0 3.8 F + Be (Dempsey90)
$\omega$ Eri 3.7 3057.0 112.5 3.7 RV constant (Merle24)
HD 70826 5.5 3900.0 25.7 3.5 G7III + (A+A) (Carquillat05)
HIP 39794 ($\epsilon$ Vol) 4.7 14.2 1.2 3.2 B6IV + B8 (Veramendi14)
HIP 30891 (V723 Mon) 5.4 59.9 1.1 3.2 Stripped giant + subgiant (ElBadry22_zoo)
@@
HIP 23900 5.3 58.3 1.4 3.1 B2V + B5-B8V (Tarasov16)
HIP 99853 (22 Vul) 2.9 249.1 3.3 3.0 Eclipsing G3I + B8V (Griffin93)
@@
While $\kappa$ Vel does not host a black hole companion, the VLTI GRAVITY observations of this system have demonstrated how the combination of RVs and high-contrast imaging may be used to effectively and efficiently vet other candidate star+black hole binaries. As described in Section §2, we selected candidates starting from the SB9 catalog (Pourbaix04) and the Gaia single-lined spectroscopic binary catalog. There are $\sim 40$ other binaries with minimum companion mass $M_{2,\rm{min}}>3\ M_{\odot}$ and projected angular separations >0.5>0.5 mas that are observable with VLTI.
@@
$\delta$ Gem: $\delta$ Gem is a hierarchical triple system (Tokovinin17). Abt65 report an orbital period 2238 days for the inner binary, $\delta$ Gem A. The binary mass function is $3.8\ M_{\odot}$ and the photometric primary is a F2 main sequence star, so the companion would have to be $\gtrsim 6\ M_{\odot}$. This orbital solution has since been questioned (Abt74), and the secondary in the inner binary has never been detected, including through an lunar occultation observation (Tremaine74). $\delta$ Gem A was also observed with the Navy Precision Optical Interferometer, but Hutter16 do not report the detection of a companion. This may be a good candidate for high-contrast imaging or interferometry. If a luminous secondary is identified, the star can be immediately rejected as a black hole candidate. If no secondary is found, additional RVs would still be needed to confirm binary orbit.
@@
HD 36881: HD 36881 is a chemically peculiar B9III star with an orbital period of 1862 d and a binary mass function of $0.85\ M_{\odot}$. Previous high-resolution spectroscopic and VLT/NACO imaging observations have not identified the companion Scholler10. We are observing this system in the current observing period 116 with VLTI GRAVITY.
@@
71 Tau: 71 Tau is $\delta$ Scuti F0IV-V star with a long orbital period (5200 d) and a large binary mass function, $f(M)=1.7\ M_{\odot}$. The companion has not been previously identified with speckle interferometry (Patience98; Mason96). Like $\delta$ Gem, the orbital ephemeris is uncertain (Abt74), and additional RVs would be needed to confirm the orbit if no companion is found with high-contrast imaging or interferometry. This star is the second brightest X-ray source in the Hyades cluster, and ultraviolet observations with the International Ultraviolet Explorer found no UV excess indicative of a hot companion (Simon97). Simon00 used Hubble Space Telescope UV spectra and identified an emission line source separated by $0\farcs 12$. Simon00 suggest that this signal is associated with an inner late-type binary companion that is chromospherically active that was undetected in previous interferometry observations that were taken near conjunction. This separation is consistent with the maximum angular separation we predict using Gaia and the SB9 orbital solution (Table 5), but Bohm07 instead suggest that the offset UV emission lines instead suggest that the emission line spectrum comes from material excited by a supernovae in the cluster.
@@
45 Leo: 45 Leo is a chemically peculiar B9IV star (Lorden83) with an uncertain long-period orbit with $f(M)=1.1\ M_{\odot}$ (Abt73). The binary has not been resolved with speckle interferometry (Hartkopf84).
@@
Table 5 also includes the 12 sources selected from Gaia DR3 that are brighter than $K<9.5$ mag, which is the magnitude limit for single-field mode observations with VLTI GRAVITY. Almost all of the targets are in the Hertzsprung gap on the color-magnitude diagram, and are likely to be binaries with an evolved red giant star and a massive main sequence star. The evolved star likely dominates the spectral energy distribution in the wavelength range of the Gaia RVS (846–870 nm) so the systems are detected SB1s.
@@
Gaia DR3 does not include the epoch RVs for the SB1 solutions, only the orbit model and covariance matrix of the posteriors, so it can be challenging to discriminate between accurate binary orbits and spurious solutions (e.g., Bashi22). For example, the Gaia SB1 solution for Gaia DR3 5824739062379517824 has a very high mass function, $f(M)=32^{+6}_{-9}\ M_{\odot}$, but the orbital period is long (>1000>1000 d) and is comparable to the baseline of the Gaia DR3 data. The photometric primary is also an early-type main sequence star with extinction-corrected $G_{\rm{BP}}-G_{\rm{RP}}<0$. Such a star will likely only have broad hydrogen lines in the Gaia RVS spectrum, so it is possible that the solution is biased by a handful of poor RV measurements. Even though the orbit is uncertain, this may still be a promising target for VLTI GRAVITY observations since a luminous companion could be identified with a small number of observations taken at different orbital phases.
@@
While there are some candidates that are suitable for interferometric or high-contrast imaging observations, this search strategy is primarily limited by the limiting magnitude of VLTI GRAVITY at small angular separations and the number of bright stars with well-characterized long period RV orbits. For a single-field off-axis observation, the limiting magnitude is $K=9.5$ mag using the ATs. New extreme adaptive optics instruments, such as GRAVITY+ (Gravity25), will allow us to push towards fainter targets ($K\lesssim 14.5$ mag). Going forward, this strategy might best be used in conjunction with RV follow-up or Gaia astrometry. For example, the catalog of spectroscopic binaries in Chini12 includes more than 80 SB1s with $K<8.5$ mag, most of which have less than 10 RVs. By combining spectroscopic monitoring of massive stars (e.g., Mahy22) with high-contrast imaging or interferometry with instruments such as GRAVITY, SHARK-VIS (Pedichini24), or CHARA (tenBrummelaar05), candidates could be vetted more efficiently than with spectroscopic follow-up alone.
@@
The number of star+compact object candidates is expected to expand dramatically with Gaia Data Release 4, which is scheduled to be released in late 2026 and will include epoch astrometry and radial velocities. Gaia BH3 (GaiaBH3), which was discovered using pre-release Gaia DR4 astrometry, has already been observed with VLTI GRAVITY to place limits on the near-infrared emission (Kervella25). Nagarajan25 predict that $\sim 30$ black holes will be detectable with Gaia DR4, and some of these long-period candidates will likely be suitable for high-contrast imaging or interferometric observations to confirm dark companions.

### v2

@@ V UVES Spectra
For each synthetic template spectrum, we use $\chi^{2}$ minimization with Nelder-Mead to determine the RV and the projected rotational velocity, $v\sin i$. We rotationally broaden the spectrum using the method described in Carvalho and Johns-Krull (2023). We do the $\chi^{2}$ minimization in two steps. First, we optimize for $v\sin i$ and the RV in $10$ Å windows around the He I lines (Figure 6), which are more sensitive to rotational broadening. Then, we fix the $v\sin i$ at the best-fit value and refine the RV with $\chi^{2}$ minimization using the full UVES CD2 wavelength range. We select the template with the lowest minimized $\chi^{2}$ as our best-fitting template.
We find the best match at $T_{\rm{eff}}=19$ kK, $\log g=3.50$ and $Z=0.5Z_{\odot}$. At Solar metallicity, we find the same best-matching effective temperature and surface gravity. The measured effective temperature and surface gravity are consistent with estimates from broadband Strömgren ubvyubvy photometry (Leone and Lanzafame, 1997). For the best fitting template, we measure $\rm{RV}=26.5$ km/s and $v\sin i=39$ km/s. This is the radial velocity we use when fitting the combined orbit model (Section §IV). The projected rotational velocity is lower than those of B-stars found in open clusters (e.g., Santos et al., 2025), but consistent with B-stars found in the Galactic field (Huang and Gies, 2008).
Figure 6 shows the UVES spectrum and the synthetic spectrum of the best-fitting B-star model and marks hydrogen Balmer, He I and metal lines. The synthetic spectrum generally reproduces the observed He I and metal lines, but less accurately matches the wings of the hydrogen Balmer lines, especially at $\lambda<4000$ Å. This is likely to represent an artifact of the continuum normalization procedure in areas of the spectrum dominated by the broad Balmer line wings. The Si II and O II lines between $4550\lesssim\lambda\lesssim 4700$ Å are also under predicted by the best matching synthetic template, which could suggest that $\kappa$ Vel is alpha-enhanced. This analysis also assumes that the spectrum is entirely dominated by the photometric primary. We evaluate this assumption in more detail in Section §VI.
The spectrum is consistent with a B2V-IV star. In addition to having no He II lines, the Si II doublet at $\lambda$4129 is clearly visible, and the C II $\lambda$4267 line is strong. The spectrum is qualitatively similar to the B2V spectroscopic standard $\beta^{2}$ Sco (Negueruela et al., 2024). The most notable differences between the best matching synthetic spectrum and the UVES spectrum of $\kappa$ Vel are the He I lines at $\lambda\lambda$4471, 4922. Figure 7 shows a zoom in on these absorption lines. In both cases, we observe an additional blended component in the blue wing of the primary absorption feature that is not represented by our synthetic spectra at any effective temperature, surface gravity, or metallicity. We consider two possibilities: (a) the additional component is actually the He I absorption line of the fainter B-star companion or (b) the line list we use with Synspec does not include this blended absorption line.
@@
While the structure in the blue wings of the He I lines at 4471Å and 4922Å are not predicted by our synthetic spectra, similar He I lines are found in other B and Be stars. For example, Negueruela et al. (2024) show that the B3V star HD 178849 has structure in the He I $\lambda$4471 line that is not apparent in stars of the same spectral class with higher rotational velocities. Similarly, the spectrum of HD 58343, a Be2V star with $v\sin i=43$ km/s also has a similar $\lambda$4471 line profile (Chauville et al., 2001) which we show in Figure 7 for comparison. The most likely explanation is that the blue component of the $\lambda$4471 comes from the forbidden transition [He I] at $\lambda$4469.96 (Daflon et al., 2007). This is is less commonly observed because it requires a high-resolution spectrum and the star to have projected rotational velocity $v\sin i\lesssim 50$ km/s (see Figure 6 of Santos et al. 2025). A forbidden [He I] line is also expected at $\lambda$4921 (Beauchamp and Wesemael, 1998), accounting for the blue wing in the He I $\lambda$4922 profile seen in Figure 7.
Figure 6: High-resolution UVES spectrum (black) taken as part of the EDIBLES survey (Cox et al., 2017). The red line shows the best fitting synthetic spectrum computed using TLUSTY model atmospheres and Synspec. Hydrogen, helium, and metal lines are marked using wavelengths from the NIST database (Kramida et al., 2024).
Figure 7: UVES spectrum of $\kappa$ Vel showing the two He I absorption lines with asymmetric blue wings not predicted by the synthetic spectra. The $\lambda$4471 line of the Be2V star HD 58343 from Chauville et al. (2001) is shown for comparison. We find that the apparent second component in the blue wings of these absorption lines is more likely from forbidden [He I] transitions rather than the stellar secondary.
VI Comparison to Evolutionary Tracks
@@
By combining the measured binary masses and the flux ratio from the VLTI/GRAVITY observations, we can estimate the evolutionary state of the binary and the spectroscopic properties of the secondary. We download precomputed evolutionary tracks (Dotter, 2016; Choi et al., 2016) from MESA Isochrones and Stellar Tracks. For a given mass pair of $M_{1}$ and $M_{2}$ and metallicity, we interpolate their tracks onto the same age grid and calculate the $K$-band flux as a function of stellar age. We can then estimate the properties of both stars and the age of the system.
@@
The measured masses are strongly dependent on the assumed parallax (Figure 5). The parallax uncertainty is large ($0.3$ and $0.48$ mas for the original and updated measurement, respectively) and the parallax could be biased by the orbital motion by as much as 22% (Section §IV). The effective temperature of the primary star measured from UVES can be used to place additional constraints on the masses of the stars. Figure 8 shows the evolution of the flux ratio of the binary system as a function of the effective temperature of the primary star assuming different parallaxes. We assume a metallicity $\rm{Fe/H}]=-0.3$ based on our template minimization results. The measured flux ratio, $\epsilon=0.285\pm 0.003$, implies dramatically different primary effective temperatures depending on the parallax and the derived primary mass. For example, assuming the original Hipparcos parallax $\varpi=6.05$ mas, the flux ratio measurement implies that the photometric primary is still about halfway through its main sequence lifetime and has an effective temperature of $T_{\rm{eff},1}>25$ kK, which is much hotter than the temperature we estimate from the UVES spectrum. Figure 8 shows that the effective temperature estimate from UVES and binary flux ratio prefer a parallax of $\varpi\approx 6.5$ mas. This is $\gtrsim 1\sigma$ larger than the original Hipparcos parallax measurement and $\gtrsim 3\sigma$ larger than the revised Hipparcos parallax measurement using the reported uncertainties, but within the broader range of parallaxes acceptable if the measurement has been biased by the binary motion.
@@
Assuming this value for the parallax, we can assess whether or not we expect to see any spectral features of the secondary in the UVES spectrum. If $\varpi\approx 6.5$ mas, the primary and secondary masses are $M_{1}=7.5\ M_{\odot}$ and $M_{2}=5.9\ M_{\odot}$, respectively. We use these masses and the orbital solution and predict that velocity difference between the primary and the secondary star at the time of the UVES observation is only $\Delta\rm{RV}=12.2$ km/s, which is not surprising given that the UVES observation was taken near conjunction. At a system age of $\sim 40$ Myr, which is when the $K$-band flux ratio matches the VLTI/GRAVITY result, both stars have $T_{\rm{eff}}\approx 19{,}000$ K, so the flux ratio is roughly the same in the $K$-band and in the wavelength range of the UVES spectrum. We take the synthetic spectra from the TLUSTY BSTAR2006 grid described in Section §V and select the closest matches to the primary and secondary star. We broaden these to match the UVES resolution and apply the expected RV shifts for each component. We also rotationally broaden the spectra of both components by 40 km/s. We combine the synthetic primary and secondary spectra, weighting them such that the flux ratio is $\epsilon=0.28$. We find that the differences between the spectrum of the photometric primary and the combined binary spectrum are at the $\lesssim 2\%$ level. It is therefore unsurprising that we do not detect any signatures of the secondary in the UVES spectrum.
@@ VII Discussion and Conclusions
$\kappa$ Vel is one of the brightest stars in the Southern sky and was first detected as a single lined spectroscopic binary by Curtis (1907). Here, we use VLTI/GRAVITY to definitively determine the nature of the secondary star for the first time. To the best of our knowledge, Daszyńska-Daszkiewicz et al. (2017) present the only previous evidence that the companion is a massive star and not a compact object. They used BRITE photometry to characterize $g$-mode oscillations and found a frequencies inconsistent with the spectral type and evolutionary state of the B2V-IV photometric primary. They instead suggest that the oscillations come from a fainter, less massive B star companion.
@@
We observed $\kappa$ Vel star as part of a broader effort to identify stars with compact object companions. Given the predicted large angular separation (Figure 1) and large companion mass (Figure 2), $\kappa$ Vel emerged as a promising candidate for VLTI/GRAVITY observations. Although were were unambiguously able to identify the luminous secondary in the first VLTI observation, we obtained a total of nine epochs, and simultaneously fit the astrometry and RVs to model the binary orbit (Figure 3). From this model, we can predict the masses of the binary components (Figure 5 and Table 4). We find that the uncertainty in the mass measurement is primarily driven by the uncertainty in the Hipparcos parallax. It is also likely that the binary motion has biased the parallax measurement by as much as $\sim 22\%$.
@@
Given its spectral type and magnitude, $\kappa$ Vel has also been targeted by studies of diffuse interstellar bands. We use the publicly available UVES spectrum from Cox et al. (2017) and B-star model atmospheres to estimate the effective temperature, surface gravity, and rotational velocity of the photometric primary (Figure 6). By combining the binary flux ratio measured from VLTI/GRAVITY and the effective temperature of the photometric primary, we can estimate the distance and binary masses (Figure 8).
@@
We also use this model to estimate how the secondary star contributes to the combined spectrum. Even at RV quadrature, when the velocity difference between the two stars is largest, the synthetic binary system is expected to differ from the photometric primary at the few-percent level (Figure 9). While it may be possible to detect the subtle spectroscopic signatures of the secondary with multi-epoch high-resolution spectra and techniques like spectral disentangling (e.g., Shenar et al., 2020; Seeburger et al., 2024), we detect the luminous secondary with the first VLTI/GRAVITY observation. Since the orbital period and ephemeris is known from the RVs, a single VLTI observation could be used to rule out a black hole with a total observing time of $\sim 1$ hour, including overheads. For comparison, the single UVES spectrum shown in Figure 6 had a total observing time of 23 min, and multiple spectra ($\gtrsim 5$) taken at different orbital phases are typically needed for spectral disentangling to be effective.
VII.1 Other VLTI Targets
@@
X Per 5.9 3900.0 6.6 20.7 Be X-ray binary pulsar (Fortin et al., 2023)
HD 45910 4.4 232.5 1.7 8.1 B2+KIII (Cowley, 1964)
$\chi$ Lup 4.0 15.3 4.1 7.2 B9.5V+ A2V (Le Bouquin et al., 2013)
$\delta$ Gem 2.6 2238.6 353.1 6.0 Uncertain RV orbit (Abt and Levy, 1974)
HD 187299 3.5 1901.0 7.4 5.9 G5I + B7V (Griffin and Griffin, 1979)
HD 50975 5.7 190.2 0.8 5.2 F8I + B2 (Sperauskas et al., 2014)
@@
HD 8556 4.9 5895.0 240.3 4.1 F0 + F5 (Fletcher, 1973)
71 Tau 4.0 5200.0 227.4 3.9 HST UV emission lines (Simon and Ayres, 2000)
HD 127208 5.1 24.6 1.0 3.8 F + Be (Dempsey et al., 1990)
$\omega$ Eri 3.7 3057.0 112.5 3.7 RV constant (Merle et al., 2024)
HD 70826 5.5 3900.0 25.7 3.5 G7III + (A+A) (Carquillat et al., 2005)
HIP 39794 ($\epsilon$ Vol) 4.7 14.2 1.2 3.2 B6IV + B8 (Veramendi and González, 2014)
HIP 30891 (V723 Mon) 5.4 59.9 1.1 3.2 Stripped giant + subgiant (El-Badry et al., 2022)
@@
HIP 23900 5.3 58.3 1.4 3.1 B2V + B5-B8V (Tarasov, 2016)
HIP 99853 (22 Vul) 2.9 249.1 3.3 3.0 Eclipsing G3I + B8V (Griffin et al., 1993)
@@
While $\kappa$ Vel does not host a black hole companion, the VLTI/GRAVITY observations of this system have demonstrated how the combination of RVs and high-contrast imaging may be used to effectively and efficiently vet other candidate star+black hole binaries. As described in Section §II, we selected candidates starting from the SB9 catalog (Pourbaix et al., 2004) and the Gaia single-lined spectroscopic binary catalog. There are $\sim 40$ other binaries with minimum companion mass $M_{2,\rm{min}}>3\ M_{\odot}$ and projected angular separations >0.5>0.5 mas that are observable with VLTI.
@@
$\delta$ Gem: $\delta$ Gem is a hierarchical triple system (Tokovinin, 2017). Abt (1965) report an orbital period 2238 days for the inner binary, $\delta$ Gem A. The binary mass function is $3.8\ M_{\odot}$ and the photometric primary is a F2 main sequence star, so the companion would have to be $\gtrsim 6\ M_{\odot}$. This orbital solution has since been questioned (Abt and Levy, 1974), and the secondary in the inner binary has never been detected, including through an lunar occultation observation (Tremaine et al., 1974). $\delta$ Gem A was also observed with the Navy Precision Optical Interferometer, but Hutter et al. (2016) do not report the detection of a companion. This may be a good candidate for high-contrast imaging or interferometry. If a luminous secondary is identified, the star can be immediately rejected as a black hole candidate. If no secondary is found, additional RVs would still be needed to confirm binary orbit.
@@
HD 36881: HD 36881 is a chemically peculiar B9III star with an orbital period of 1862 d and a binary mass function of $0.85\ M_{\odot}$. Previous high-resolution spectroscopic and VLT/NACO imaging observations have not identified the companion (Schöller et al., 2010). We are observing this system in the current observing period 116 with VLTI/GRAVITY.
@@
71 Tau: 71 Tau is $\delta$ Scuti F0IV-V star with a long orbital period (5200 d) and a large binary mass function, $f(M)=1.7\ M_{\odot}$. The companion has not been previously identified with speckle interferometry (Patience et al., 1998; Mason, 1996). Like $\delta$ Gem, the orbital ephemeris is uncertain (Abt and Levy, 1974), and additional RVs would be needed to confirm the orbit if no companion is found with high-contrast imaging or interferometry. This star is the second brightest X-ray source in the Hyades cluster, and ultraviolet observations with the International Ultraviolet Explorer found no UV excess indicative of a hot companion (Simon and Landsman, 1997). Simon and Ayres (2000) used Hubble Space Telescope UV spectra and identified an emission line source separated by $0\farcs 12$. Simon and Ayres (2000) suggest that this signal is associated with an inner late-type binary companion that is chromospherically active that was undetected in previous interferometry observations that were taken near conjunction. This separation is consistent with the maximum angular separation we predict using Gaia and the SB9 orbital solution (Table 5), but Böhm-Vitense (2007) instead suggest that the offset UV emission lines come from material excited by a supernovae in the cluster.
@@
45 Leo: 45 Leo is a chemically peculiar B9IV star (Loden, 1983) with an uncertain long-period orbit with $f(M)=1.1\ M_{\odot}$ (Abt and Snowden, 1973). The binary has not been resolved with speckle interferometry (Hartkopf and McAlister, 1984).
@@
Table 5 also includes the 12 sources selected from Gaia DR3 that are brighter than $K<9.5$ mag, which is the magnitude limit for single-field mode observations with VLTI/GRAVITY. Almost all of the targets are in the Hertzsprung gap on the color-magnitude diagram, and are likely to be binaries with an evolved red giant star and a massive main sequence star. The evolved star likely dominates the spectral energy distribution in the wavelength range of the Gaia RVS (846–870 nm) so the systems are detected SB1s.
@@
Gaia DR3 does not include the epoch RVs for the SB1 solutions, only the orbit model and covariance matrix of the posteriors, so it can be challenging to discriminate between accurate binary orbits and spurious solutions (e.g., Bashi et al., 2022). For example, the Gaia SB1 solution for Gaia DR3 5824739062379517824 has a very high mass function, $f(M)=32^{+6}_{-9}\ M_{\odot}$, but the orbital period is long (>1000>1000 d) and is comparable to the baseline of the Gaia DR3 data. The photometric primary is also an early-type main sequence star with extinction-corrected $G_{\rm{BP}}-G_{\rm{RP}}<0$. Such a star will likely only have broad hydrogen lines in the Gaia RVS spectrum, so it is possible that the solution is biased by a handful of poor RV measurements. Even though the orbit is uncertain, this may still be a promising target for VLTI/GRAVITY observations since a luminous companion could be identified with a small number of observations taken at different orbital phases.
@@
While there are some candidates that are suitable for interferometric or high-contrast imaging observations, this search strategy is primarily limited by the limiting magnitude of VLTI/GRAVITY at small angular separations and the number of bright stars with well-characterized long period RV orbits. For a single-field off-axis observation, the limiting magnitude is $K=9.5$ mag using the ATs. New extreme adaptive optics instruments, such as GRAVITY+ (GRAVITY+ Collaboration, 2025), will allow us to push towards fainter targets ($K\lesssim 14.5$ mag). Going forward, this strategy might best be used in conjunction with RV follow-up or Gaia astrometry. For example, the catalog of spectroscopic binaries in Chini et al. (2012) includes more than 80 SB1s with $K<8.5$ mag, most of which have less than 10 RVs. By combining spectroscopic monitoring of massive stars (e.g., Mahy et al., 2022) with high-contrast imaging or interferometry with instruments such as GRAVITY, SHARK-VIS (Pedichini et al., 2024), or CHARA (ten Brummelaar et al., 2005), candidates could be vetted more efficiently than with spectroscopic follow-up alone.
@@
The number of star+compact object candidates is expected to expand dramatically with Gaia Data Release 4, which is scheduled to be released in late 2026 and will include epoch astrometry and radial velocities. Gaia BH3 (Gaia Collaboration et al., 2024), which was discovered using pre-release Gaia DR4 astrometry, has already been observed with VLTI/GRAVITY to place limits on the near-infrared emission (Kervella et al., 2025). Nagarajan et al. (2025) predict that $\sim 30$ black holes will be detectable with Gaia DR4, and some of these long-period candidates will likely be suitable for high-contrast imaging or interferometric observations to confirm dark companions.